QUESTION 8 When the pavement is dry, a car with a mass of 950 kg can drive as fast as 70 km/h on a curve without skidding. When it rains, the coefficient of static friction is 2/3 of the coefficient of static friction under dry conditions. How fast can a 1,100kg car drive on the same curve while raining? 47km/h O 82 km/h 57 km/h 54 km/h ● 66km/h
Answer #1

from banking angle

tan theta = us = V^2/Rg

V = 70 kmph = 70 * 5/18 = 19.44 m/s

radius of the circle R = v^2/(us* g)

R = 19.44^2/(us * 9.81)

R = 38.52/us m


now v^2 = R g tan theta

V^2 = (38.52/us * 9.81* 2/3 * us)

V = 15.87 m/s

V = 15.87* 18/5 = 57.14 kmph

option three it is